3. The rate of decomposition of N2O5 in the reaction 2 N2O5(g)  4 NO2(g) + 5 O2(g) at a particular instant is 4.2 x 10-7 M/s, what is the rate of appearance of NO2?
Answer #1

The rate of decomposition of N2O5 in the reaction 2 N2O5(g)  4 NO2(g) + 5 O2(g) at a particular instant is 4.2 x 10-7 M/s, what is the rate of appearance of NO2?

we have :

2 N2O5(g) ==> 4 NO2(g) + 5 O2(g)

Rate of decomposition of N2O5 = -d[N2O5] /dt = 4.2 x 10-7 M/s

we have, Rate of reaction = - 1/2*d[N2O5] /dt = 1/4*d[NO2] /dt = 2.1 x 10-7 M/s

thus rate of appearance of NO2 : d[NO2] /dt = 4*2.1 x 10-7 M/s = 8.4 x 10-7 M/s